1. Fill in all the remaining blanks in the table below and fill out the appropriate formula each.
Residual maturity

Residual Maturity Interest rate of Zero coupon bond Price of Zero Coupon Bond one-year intrinsic forward interest rate par yield
1 8.00%
2 7.00%
3 6.00%
Answer #1
Residual Maturity Interest rate of Zero coupon bond Price of Zero Coupon Bond one-year intrinsic forward interest rate par yield
1 8.00% =1000/(1+8%)=925.26 =((1+7%)^2/(1+8%)^1)-1=6.01% 8%
2 7.00% =1000/(1+7%)^2=873.4 =((1+6%)^3/(1+7%)^2)-1=4.03% 7%
3 6.00% =1000/(1+6%)^3=839.6 6%